# Abelian congruences and similarity in varieties with a weak difference term

### v1

@@ 8. Examples
    $K_{\ell}(a,b)=\left\{\begin{array}[]{cl}b&\mbox{if $a\in C_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\not\in C_{\ell}$ and $b\in C_{k}$}.\end{array}\right.$
Let $\mathscr{F}$ be the set of all the operations $F_{(\ell,i)}^{g}$, let $\mathscr{G}$ be the set of all the operations $G_{\ell}$ and $G_{\ell}^{\prime}$, let =$\mathscr{H}=\{H_{s}\,:\,\text{$s\in S$}\}$, and let $\mathscr{K}=\{K_{\ell}\,:\,\text{$1\leq\ell\leq m$}\}$. Finally, let
@@
    $\Delta_{\mu,\alpha}$ is the set of all pairs $((a,b),(a^{\prime},b^{\prime}))$ such that for some $\ell\leq m$ and some $i,j\leq n_{\ell}$, $\{a,b\}\subseteq V_{(\ell,i)}$ and $\{a^{\prime},b^{\prime}\}\subseteq V_{(\ell,j)}$ and
@@
(1) follows from Lemma 8.3. It is easy to check that $\mu$ and $\alpha$ are congruences of ${\mathbf{A}}$. Let $\theta\in\operatorname{Con}({\mathbf{A}})$ satisfy $0\neq\theta\subseteq\mu$. (We want to show $\theta=\mu$.) We can choose $(\ell,i)\in S$ so that $\theta\cap(V_{(\ell,i)})^{2}$ contains a nontrivial pair $(c,c^{\prime})$, i.e., with $c\neq c^{\prime}$. Let $a=c-c^{\prime}$ calculated in $V_{(\ell,i)}$. Then $a\neq 0_{(\ell,i)}$ and $(a,0_{(\ell,i)})\in\theta$ because of $d(x,y,z)$. Now let $j\leq n_{\ell}$ and $b\in V_{(\ell,j)}$ with $b\neq 0_{(\ell,j)}$. Let $a^{\prime}=\sigma_{(\ell,i)}(a)$. Then $a^{\prime}\neq 0_{(\ell,0)}$. Hence there exists an $\mathbb{F}$-linear map $g:V_{(\ell,0)}\rightarrow V_{(n,j)}$ with $g(a^{\prime})=b$. Then $F_{(\ell,j)}^{g}(a)=b$ while $F_{(\ell,j)}^{g}(0_{(\ell,i)})=0_{(\ell,j)}$, proving $(b,0_{(\ell,j)})\in\theta$. As $j$ and $b$ were arbitrary, this proves $\mu\cap(C_{\ell})^{2}\subseteq\theta$.
If $\ell>0$, then we can use the operation $G_{\ell}^{\prime}$ to show that $\theta\cap(V_{(0,0)})^{2}$ contains a nontrivial pair. Then the argument in the previous paragraph gives $\mu\cap(C_{0})^{2}\subseteq\theta$. Now let $k\in\{1,\ldots,m\}$ be such that some $\mu$-class $V_{(k,j)}$ contained in $C_{k}$ has more than one element. Then we can use the operation $G_{k}$ to show that $\theta\cap(V_{(k,0)})^{2}$ contains a nontrivial pair. Again by the argument in the previous paragraph, we get $\mu\cap(C_{k})^{2}\subseteq\theta$. This finishes the proof that $\theta=\mu$; hence $0_{A}\prec\mu$.
Next let $(c,c^{\prime})$ be any nontrivial pair in $A^{2}$ with $(c,c^{\prime})\not\in\mu$. Say $c\in V_{s}$ and $c^{\prime}\in V_{t}$ with $s\neq t$. Then the polynomial $f(x)=H_{s}(x,c)$ sends $(c,c^{\prime})$ to $(1,0_{(0,0)})$, so $\operatorname{Cg}(c,c^{\prime})\cap\mu\neq 0$. As we have shown $0_{A}\prec\mu$, this proves $\mu$ is the monolith of ${\mathbf{A}}$. Combining the above argument with Corollary 3.2, we also get that for all $(a,b)\in\mu$ and all $(c,c^{\prime})\in A^{2}$ with $c\neq c^{\prime}$, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ wth $f(c)=a$ and $f(c^{\prime})=b$. This proves that $\mu$ is the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
@@
proving $R\subseteq\operatorname{Sg}^{{\mathbf{A}}^{2\times 2}}(X(\mu,\alpha))$. These remarks prove $R=M(\mu,\alpha)$. Since every matrix in $R$ has the property that the first column is in $0_{A}$ if and only if the second column is in $0_{A}$, we have proved that $C(\alpha,\mu;0)$ holds and therefore $\mu$ is abelian and $(0:\mu)\geq\alpha$. To prove $(0:\mu)=\alpha$, suppose $\theta\in\operatorname{Con}({\mathbf{A}})$ and $\theta>\alpha$. Pick $(a,b)\in\theta\setminus\alpha$, say with $a\in C_{\ell}$ and $b\in C_{k}$ with $k\neq\ell$. Using the operation $K_{\ell}$ and the pair $(0_{(0,0)},1)\in\mu\setminus 0_{A}$, we have
@@
Recall that $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$. Let $\Delta=\Delta_{\mu,\alpha}$ and $D({\mathbf{A}})={\mathbf{A}}(\mu)/\Delta$. Observe that every element of $D({\mathbf{A}})$ can be uniquely written in the form $(a,0_{(\ell,0)})/\Delta$ for some $\ell\leq m$ and $a\in V_{(\ell,0)}$. Let $h:D({\mathbf{A}})\rightarrow B$ be the bijection sending $(a,0_{(\ell,0)})/\Delta\mapsto a$. Then by construction, for every $(\ell,i)\in S$ we have
@@
=$=\{(a,0_{(\ell,i)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$
=$=\{(\sigma_{(\ell,i)}(a),0_{(\ell,0)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$ Claim 8.4.1(4)
=$=\{(a^{\prime},0_{(\ell,0)})/\Delta\,:\,\text{$a^{\prime}\in W_{(\ell,i)}$}\}$ as $\sigma_{(\ell,i)}(V_{(\ell,i)})=W_{(\ell,i)}$.
@@
Finally, let $T$ be the transversal for $\mu$ given by $T=\{0_{s}\,:\,\text{$s\in S$}\}$, and let $\mathbb{F}^{\prime}$ be the division ring associated to ${\mathbf{A}},\mu,T$. We will verify that $\mathbb{F}^{\prime}\cong\mathbb{F}$. For each $\zeta\in\mathbb{F}$ and $s\in S$ let $\zeta_{s}:V_{s}\rightarrow V_{s}$ be the “scalar multiplication by $\zeta$ in $V_{s}$” map and let $\Theta(\zeta)=(\zeta_{s})_{s\in S}$. Every $n$-ary basic operation of ${\mathbf{A}}$, when restricted to an $n$-tuple of $\mu$-classes, is $\mathbb{F}$-affine, and this property is inherited by polynomial operations. It follows that $\Theta(\zeta)\in\mathbb{F}^{\prime}$. To prove $\operatorname{ran}(\Theta)=\mathbb{F}^{\prime}$, let $\lambda=(\lambda_{s})_{s\in S}\in\mathbb{F}^{\prime}$. Then
@@
Fix $\ell\leq m$ with $\dim(V_{(\ell,0)})>0$. Since every $\mathbb{F}$-linear map $V_{(\ell,0)}\rightarrow V_{(\ell,0)}$ is a restriction to $V_{(\ell,0)}$ of some unary polynomial, ($\ast$8.4) implies that $\lambda_{(\ell,0)}$ commutes with every $r\in\operatorname{End}_{\mathbb{F}}(V_{(\ell,0)})$. Hence there exists $\zeta=\zeta^{(\ell)}\in\mathbb{F}$ such that $\lambda_{(\ell,0)}=\zeta_{(\ell,0)}$. If $\ell\neq 0$, then recall the nonconstant $\mathbb{F}$-linear map $g_{\ell}:V_{(0,0)}\rightarrow V_{(\ell,0)}$ used to define the basic operation $G_{\ell}$. Then $g_{\ell}=G_{\ell}{\upharpoonright}_{V_{(0,0)}}$, so ($\ast$8.4) gives $\lambda_{(\ell,0)}\circ g_{\ell}=g_{\ell}\circ\lambda_{(0,0)}$. Let $\zeta=\zeta^{(0,0)}$ and $\zeta^{\prime}=\zeta^{(\ell,0)}$. Choose $a\in V_{(0,0)}$ with $b:=g_{\ell}(a)\neq 0_{(\ell,0)}$. Then
$\zeta\cdot b=\zeta\cdot g_{\ell}(a)$ $=g_{\ell}(\zeta\cdot a)$ $\mathbb{F}$-linearity of $g_{\ell}$
$=g_{\ell}(\zeta_{(0,0)}(a))$ definition of $\zeta_{(0,0)}$
$=g_{\ell}(\lambda_{(0,0)}(a))$
$=\lambda_{(\ell,0)}(g_{\ell}(a))$ by ($\ast$8.4)
$=\zeta^{\prime}\cdot b.$
Since $b\neq 0_{(\ell,0)}$, we get $\zeta=\zeta^{\prime}$, i.e., $\zeta^{(0,0)}=\zeta^{(\ell,0)}$. A similar argument using operations in $\mathscr{F}$ gives $\zeta^{(\ell,i)}=\zeta^{(\ell,0)}$ whenever $\ell\leq m$ and $\dim(V_{(\ell,i)})>0$. Hence there exists $\zeta\in\mathbb{F}$ such that $\zeta^{s}=\zeta$ for all $s\in S$ with $\dim(V_{s})>0$. It follows that $\lambda=\Theta(\zeta)$, so $\Theta(\mathbb{F})=\mathbb{F}^{\prime}$. The reader can verify that $\Theta:\mathbb{F}\cong\mathbb{F}^{\prime}$.
9. Conclusion and questions
In this paper we have extended Freese’s analysis of abelian minimal congruences, and his similarity relation between subdirectly irreducible algebras, from the congruence modular setting to the setting of varieties with a weak difference term. We also established a connection between similarity and certain constructs which we have called “proper bridges.” In the companion paper [33], we will compare proper bridges to Zhuk’s original bridges (which are superficially weaker) and explain some related elements of Zhuk’s work in terms of centrality and similarity. In a third paper [32], we will unite the analyses of “rectangular critical relations” by Zhuk [37, Lemma 8.21] (in the idempotent locally finite Taylor case) and by Kearnes and Szendrei [20] (in the congruence modular case), showing in particular the role of similarity, associated finite fields, and linear equations.
Problem 1.
Because of their connection to the Constraint Satisfaction Dichotomy Theorem, finite subdirectly irreducible algebras with abelian monoliths in an idempotent locally finite Taylor variety are of particular interest. We have seen one result (Lemma 5.9) showing how the assumption of idempotency can simplify our analysis. In what other ways does idempotency affect the analysis? For example:
Problem 2.
Suppose ${\mathbf{A}}$ is a finite idempotent subdirectly irreducible algebra in a Taylor variety with abelian monolith $\mu$, $\alpha=(0:\mu)$, and $C$ is an $\alpha$-class. If $q$ is the maximal size of a $\mu$-class in $C$, does it follow that every $\mu$-class in $C$ has size $q$ or 1?

### v2

@@ 8. Conclusion and questions
    $K_{\ell}(a,b)=\left\{\begin{array}[]{cl}b&\mbox{if $a\in E_{\ell}$}\\ 0_{(k,0)}&\mbox{if $a\not\in E_{\ell}$ and $b\in E_{k}$}.\end{array}\right$.
Let $\mathscr{F}$ be the set of all the operations $F_{(\ell,i)}^{g}$, let $\mathscr{G}$ be the set of all the operations $G_{\ell}$ and $G_{\ell}^{\prime}$, let $\mathscr{H}=\{H_{s}\,:\,\text{$s\in S$}\}$, and let $\mathscr{K}=\{K_{\ell}\,:\,\text{$1\leq\ell\leq m$}\}$. Finally, let
@@
    $\Delta_{\mu,\alpha}$ is the set of all pairs $((a,b),(a^{\prime},b^{\prime}))$ such that for some $\ell\leq m$ and some $i,j\leq n_{\ell}$, $\{a,b\}\subseteq V_{(\ell,i)}$ while $\{a^{\prime},b^{\prime}\}\subseteq V_{(\ell,j)}$ and
@@
(1) follows from Lemma A2.3. It is easy to check that $\mu$ and $\alpha$ are congruences of $\mathbf{A}$}. Let $\theta\in\operatorname{Con}({\mathbf{A}})$ satisfy $0\neq\theta\subseteq\mu$. (We want to show $\theta=\mu$.) We can choose $(\ell,i)\in S$ so that $\theta\cap(V_{(\ell,i)})^{2}$ contains a pair $(c,c^{\prime})$ with $c\neq c^{\prime}$. Let $a=c-c^{\prime}$ calculated in $V_{(\ell,i)}$. Then $a\neq 0_{(\ell,i)}$ and $(a,0_{(\ell,i)})\in\theta$ because of $d(x,y,z)$. Now let $j\leq n_{\ell}$ and $b\in V_{(\ell,j)}$ with $b\neq 0_{(\ell,j)}$. Let $a^{\prime}=\sigma_{(\ell,i)}(a)$. Then $a^{\prime}\neq 0_{(\ell,0)}$. Hence there exists an $\mathbb{F}$-linear map $g:V_{(\ell,0)}\rightarrow V_{(n,j)}$ with $g(a^{\prime})=b$. Then $F_{(\ell,j)}^{g}(a)=b$ while $F_{(\ell,j)}^{g}(0_{(\ell,i)})=0_{(\ell,j)}$, proving $(b,0_{(\ell,j)})\in\theta$. As $j$ and $b$ were arbitrary, this proves $\mu\cap(E_{\ell})^{2}\subseteq\theta$.
If $\ell>0$, we can use the operation $G_{\ell}^{\prime}$ to show that $\theta\cap(V_{(0,0)})^{2}$ contains a nontrivial pair. Then the argument in the previous paragraph gives $\mu\cap(E_{0})^{2}\subseteq\theta$. Now let $k\in\{1,\ldots,m\}$ be such that some $\mu$-class $V_{(k,j)}$ contained in $E_{k}$ has more than one element. Then we can use the operation $G_{k}$ to show that $\theta\cap(V_{(k,0)})^{2}$ contains a nontrivial pair. Again by the argument in the previous paragraph, we get $\mu\cap(E_{k})^{2}\subseteq\theta$. This finishes the proof that $\theta=\mu$; hence $0_{A}\prec\mu$.
Next let $(c,c^{\prime})$ be any nontrivial pair in $A^{2}$ with $(c,c^{\prime})\not\in\mu$. Say $c\in V_{s}$ and $c^{\prime}\in V_{t}$ with $s\neq t$. Then the polynomial $f(x)=H_{s}(x,c)$ sends $(c,c^{\prime})$ to $(1,0_{(0,0)})$, so $\operatorname{Cg}(c,c^{\prime})\cap\mu\neq 0$. As we have shown $0_{A}\prec\mu$, this proves $\mu$ is the monolith of $\mathbf{A}$}. Combining the above argument with Corollary 3.2, we also get that for all $(a,b)\in\mu$ and all $(c,c^{\prime})\in A^{2}$ with $c\neq c^{\prime}$, there exists $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(c)=a$ and $f(c^{\prime})=b$. This proves that $\mu$ is the unique minimal reflexive subuniverse of ${\mathbf{A}}^{2}$ properly containing $0_{A}$.
@@
proving $R\subseteq\operatorname{Sg}^{{\mathbf{A}}^{2\times 2}}(X(\mu,\alpha))$. These remarks prove $R=M(\mu,\alpha)$. Since every matrix in $R$ has the property that the first column is in $0_{A}$ if and only if the second column is in $0_{A}$, we have proved that $C(\alpha,\mu;0)$ holds and therefore $\mu$ is abelian and $(0:\mu)\geq\alpha$. To prove $(0:\mu)=\alpha$, suppose $\theta\in\operatorname{Con}({\mathbf{A}})$ and $\theta>\alpha$. Pick $(a,b)\in\theta\setminus\alpha$, say with $a\in E_{\ell}$ and $b\in E_{k}$ with $k\neq\ell$. Using the operation $K_{\ell}$ and the pair $(0_{(0,0)},1)\in\mu\setminus 0_{A}$, we have
@@
Recall that $B=V_{(0,0)}\cup\cdots\cup V_{(m,0)}$. Let $\Delta=\Delta_{\mu,\alpha}$, so $D({\mathbf{A}},\mu)={\mathbf{A}}(\mu)/\Delta$. Observe that every element of $D({\mathbf{A}},\mu)$ can be uniquely written in the form $(a,0_{(\ell,0)})/\Delta$ for some $\ell\leq m$ and $a\in V_{(\ell,0)}$. Let $h:D({\mathbf{A}},\mu)\rightarrow B$ be the bijection sending $(a,0_{(\ell,0)})/\Delta\mapsto a$. Then by construction, for every $(\ell,i)\in S$ we have
@@
$=\{(a,0_{(\ell,i)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$
$=\{(\sigma_{(\ell,i)}(a),0_{(\ell,0)})/\Delta\,:\,\text{$a\in V_{(\ell,i)}$}\}$ Claim A2.4(4)
$=\{(a^{\prime},0_{(\ell,0)})/\Delta\,:\,\text{$a^{\prime}\in W_{(\ell,i)}$}\}$ as $\sigma_{(\ell,i)}(V_{(\ell,i)})=W_{(\ell,i)}$.
@@
Finally, let $T$ be the transversal for $\mu$ given by $T=\{0_{s}\,:\,\text{$s\in S$}\}$, and let $\mathbb{D}$ be the division ring associated to $\mu,T$ by Freese’s construction as described in Appendix 1. We will verify that $\mathbb{F}\cong\mathbb{D}$, which will imply $\mathbb{F}\cong\mathbb{F}_{\mu}$ by Proposition A1.1.
As in the proof of Proposition A1.1, let $V$ be the group $\prod_{s\in S}\operatorname{Grp}_{{\mathbf{A}}}(\mu,0_{s})$, let $\mathbf{P}$} be the ring $\prod_{s\in S}\mathbf{End}(\operatorname{Grp}_{{\mathbf{A}}}(\mu,0_{s}))$, let $\Psi:{\mathbf{P}}\rightarrow\mathbf{End}(V)$ be the obvious (coordinatewise) ring embedding, and recall that $\mathbb{D}$ is the subring of $\mathbf{End}(V)$ whose universe is
$\{\Psi((\lambda_{s})_{s\in S})\,:\,\text{$\lambda_{s}\circ f{\upharpoonright}_{V_{t}}=f\circ\lambda_{t}$ for all $f\in\operatorname{Pol}_{1}({\mathbf{A}})$ with $f(0_{t})=0_{s}$}\}.$
For each $\zeta\in\mathbb{F}$ and $s\in S$ let $\zeta_{s}:V_{s}\rightarrow V_{s}$ be the “scalar multiplication by $\zeta$ in $V_{s}$” map and let $\Theta(\zeta)=(\zeta_{s})_{s\in S}$. Clearly $(\zeta_{s})_{s\in S}\in P$. Every $n$-ary basic operation of $\mathbf{A}$}, when restricted to an $n$-tuple of $\mu$-classes, is $\mathbb{F}$-affine, and this property is inherited by polynomial operations. It follows that $(\Psi\circ\Theta)(\zeta)\in\mathbb{D}$. It is not hard to check that $\Psi\circ\Theta$ is a ring embedding of $\mathbb{F}$ into $\mathbb{D}$.
To prove $\operatorname{ran}(\Psi\circ\Theta)=\mathbb{D}$, fix $\lambda\in\mathbb{D}$ and choose $(\lambda_{s})_{s\in S}\in P$ with $\Psi((\lambda_{s})_{s\in S})=\lambda$. Thus
@@
[…]
$\zeta^{t}\cdot b=\zeta^{t}\cdot g_{\ell}(a)$ $=g_{\ell}(\zeta^{t}\cdot a)$ $\mathbb{F}$-linearity of $g_{\ell}$
$=g_{\ell}(\lambda_{t}(a))$
[…]
$=\zeta^{s}\cdot b.$
Since $b\neq 0_{s}$, we get $\zeta^{s}=\zeta^{t}$. A similar argument using operations in $\mathscr{F}$ gives $\zeta^{(\ell,i)}=\zeta^{(\ell,0)}$ whenever $\ell\leq m$ and $\dim(V_{(\ell,i)})>0$. This completes the proof that $\mathbb{F}_{u}\cong\mathbb{F}$.